# Open-Vocabulary SAM: Segment and Recognize Twenty-thousand Classes Interactively

### v1

@@ 4.2 Ablation Studies and Analysis
| Setting | IoU | Acc | ||
| SAM2CLIP (baseline) | 78.1 | 54.2 | 27.6 | 33.2 |
| + Cls Token | 81.3 | 79.3 | 29.8 | 34.5 |
| + Cls Token & CLIP MLP fusion | 80.9 | 78.9 | 29.0 | 33.9 |
@@
Detailed Design on CLIP2SAM. In Tab. 6, we present extensive design for the CLIP2SAM module. We compare two designs: a simple classification token with cross attention (Cls Token) and a combination of this token with mask pooled CLIP feature (CLS Token & CLIP MLP fusion). These designs work better than the combined baseline shown in the first row. Nonetheless, due to resolution constraints, these variants cannot handle small objects well, as shown in Fig. 4. In contrast, our design that includes a light FPN improves the performance considerably.
| Setting | IoU | Acc | #GFlop(G) | #Params (M) |
@@
Ablation on Different CLIP Backbones. In Tab. 7, we explore the effect of frozen CLIP visual backbone. We do not add the CLIP2SAM module. Motivated by recent works [28, 72, 70, 64], CNN-based CLIPs encapsulate more structural information, which is good for our goal since we have location-sensitive visual prompts as the input. Thus, we avoid naïve ViT design for SAM2CLIP but adopt CNN-based CLIPs. As shown in the table, we find ConvNext large achieves the best performance.
@@ 5 Conclusion
We present Open Vocabulary SAM, a SAM-inspired method for interactive segmentation and recognition. Unlike previous open-vocabulary detection and segmentation methods, our method explores interactive open-vocabulary segmentation for the first time. Given the user’s inputs, such as boxes or points, the proposed approach can interactively segment and label each visual prompt. Compared with the combined baselines and various visual adapters, our proposed CLIP2SAM and SAM2CLIP are both efficient and effective in various settings. Our open vocabulary segmentation is compatible with different detectors, including open-vocabulary detectors and close-set detectors. With more data, our model plays a similar role as SAM, offering an effective annotation tool for both segmentation and instance labeling. In particular, our method can perform large vocabulary segmentation and recognition over 22K classes. We hope our research can provide a solid baseline for combining the strengths of different forms of vision foundation models.
Broader Impact and Limitations. Our study advances interactive segmentation and recognition by combining CLIP and SAM into one framework. This integration holds substantial potential to enhance annotation and image editing applications. A current limitation is the lack of exploration into mask prompts for interactive tasks, which we plan to address in future work.
@@
Overview. In the supplementary materials, we provide more details and results to support our main paper. The contents are presented as follows: we first present more details and discussions of our method in Sec. A. Then, we report more results in Sec. B. Finally, we present a video demo for method introduction and discuss future work in Sec. C (see introduction.mp4 and tools_demo.mov).
| Detector | Method | COCO | LVIS | |||||||
| #Params (M) | FLOPs (G) | |||||||||
| Box Prompt (box AP) | 42.2 | 52.3 | 44.9 | 45.3 | 45.8 | 48.4 | 46.7 | - | - | |
| Detic [86] (Swin-Base) | SAM | 35.2 | 43.1 | 37.2 | 42.6 | 43.3 | 44.2 | 43.5 | 641 | 3,001 |
| Open-Vocabulary SAM | 36.3 | 43.6 | 38.2 | 42 | 42.5 | 42.9 | 42.6 | 304 | 1,180 | |
| Box Prompt (box AP) | 36.8 | 22.5 | 33.1 | 36.7 | 36.4 | 38.7 | 37.3 | - | - | |
| CLIPSelf [64] (Swin-Base) | SAM | 37.0 | 23.6 | 33.5 | 39 | 38.9 | 40.8 | 39.7 | 641 | 3,001 |
| Open-Vocabulary SAM | 35.7 | 21.6 | 32.1 | 36.8 | 36.8 | 37.7 | 37.2 | 304 | 1,180 | |
| Box Prompt (box AP) | 57.3 | 65.1 | 59.3 | 39.1 | 51.3 | 57.2 | 51.5 | - | - | |
| ViTDet [39] (Huge) | SAM | 44.8 | 50.6 | 46.3 | 34.6 | 45.4 | 47.8 | 44.5 | 641 | 3,001 |
| Open-Vocabulary SAM | 47.5 | 52.6 | 48.8 | 34.5 | 44.9 | 47.3 | 44.0 | 304 | 1,180 | |
Appendix A Method Discussion
Comparison with Recent Joint SAM and CLIP Models. Several recent works [29, 79, 56] also explore joint segmentation and recognition as one system. Recognize Anything [79] adopts the tagging-based model. However, the focus of that work is to build large-scale tagging datasets via automatic text semantic parsing. Moreover, it cannot perform interactive segmentation. SAM-CLIP [56] integrates multi-task learning, continual learning techniques, and teacher-student distillation, where its goal is to build semantic segmentation models. On the other hand, our Open-Vocabulary SAM is a variant of SAM, which can support interactive segmentation and mask classification with flexible user inputs. Therefore, SAM-CLIP is orthogonal to the proposed Open-Vocabulary SAM. Meanwhile, Semantic-SAM [29] extends Mask-DINO [30] with fine-grained and interactive segmentation. It contains a decoupled transformer decoder to generate entity proposals, making it a universal and multi-task model. In contrast, our model only contains a lightweight decoder (SAM decoder) and CLIP2SAM prediction. Besides, our work mainly explores the open-vocabulary ability of a frozen CLIP model to enhance the SAM. As a result, our method only requires training partial parameters and fuses the knowledge embodied in SAM and CLIP.
@@ Appendix A Method Discussion
Comparison with Open-Vocabulary Methods. Our work is orthogonal to previous open-vocabulary methods. Specifically, we aim to build a SAM-like model with interactive prompts, such as points and boxes. Previous open-vocabulary methods need to generate region proposals or mask proposals to recall all possible objects in the scene. Following the same spirit of SAM, our model mainly takes these proposals for segmentation and recognition. Thus, we can deploy our model on various OV-detectors [86, 64] to achieve instance segmentation or class-agnostic detectors to achieve joint segmentation and recognition.
Training and Inference Details. In Fig. 6, we present a more detailed visualization of both training and inference of our Open-Vocabulary SAM. As shown in the figure, only three components are learned during the training, including SAM2CLIP (129M), CLIP2SAM (3.8M), and SAM decoder (4.0M). The remaining parameters are frozen. The total parameters are 304M. After SAM2CLIP distillation, the heavy SAM encoder (637M) is dropped when fine-tuning our model on the detection and classification datasets, which speeds up the CLIP2SAM process.
@@
The CLIP2SAM module is co-trained using a diverse mixture of datasets, encompassing detection, segmentation, and classification tasks. Specifically, for segmentation, the training integrates both mask and classification labels, while for detection and classification tasks, the focus is solely on training the classification head. The classification dataset, notable for its extensive range of class categories, coupled with the inclusion of several recent works that feature a large vocabulary, significantly enhances the model’s capabilities. After this extensive co-training process, our model exhibits the remarkable ability to recognize and segment over twenty thousand class categories, all within a single, unified framework.
Appendix B More Experimental Results
Comparative Results on More Detectors. In Tab. 8, we compare the Open-Vocabulary SAM with the original SAM (ViT-H) across various detectors. Our method uses the scores and labels generated by the corresponding detectors and bounding boxes as prompts for mask generation. Notably, with ViTDet (Huge), a strong detector, our Open-Vocabulary SAM achieves comparable or superior segmentation to the original SAM. With less robust detectors, Open-Vocabulary SAM segmentation ability slightly decreases. It is noteworthy that the Open-Vocabulary SAM is more efficient, requiring fewer parameters and less computational resources than the SAM (ViT-H) model.
More Comparisons with SAM Models. In Tab. 9, we compare our method with several SAM models, focusing on open-vocabulary recognition. We generate masks using ground truth bounding boxes from the COCO dataset [41] and evaluate them using the per-instance IoU metric (1-IoU), as per Semantic-SAM [29]. Our Open-Vocabulary SAM stands out for its ability to recognize a broad spectrum of categories in an open-vocabulary setting, in addition to effective mask generation.
Scaling Up With More Datasets. Our Open-Vocabulary SAM, scaled with large-scale datasets, demonstrates impressive zero-shot classification on the COCO validation set, detailed in Tab.10. Using ImageNet-21k (I-21k)[11], a dataset with 19,167 categories and image-level supervision, the model achieves a 44.5% accuracy in zero-shot instance-level classification. This underscores the efficacy of training with readily available image-level annotations. In addition, incorporating large-vocabulary detection datasets like LVIS [18], Objects365 [53], and V3Det [57], which offer instance-level annotations, significantly enhances classification accuracy. For a more comprehensive understanding of the model’s open-vocabulary classification capabilities, we direct readers to watch our demo video on our project page https://www.mmlab-ntu.com/project/ovsam.
| Datasets | Accuracy (COCO) | #vocaulary | #images |
| LVIS | 83.1 | 1,203 | 99K |
| V3Det | 78.7 | 13,204 | 183K |
| I-21k | 44.5 | 19,167 | 13M |
| V3Det + LVIS | 82.7 | 13,844 | 282K |
| V3Det + LVIS + I-21k | 83.3 | 25,898 | 13M |
| V3Det + LVIS + I-21k + Object365 | 83.0 | 25,970 | 15M |
Appendix C Failure Cases, Demo and Future Work
Failure Cases. In Fig. 8, we show two instances where our Open-Vocabulary SAM falls short. The first case, depicted on the left, involves class labels that are subtly distinct and challenging to differentiate, even for humans. The second scenario, shown on the right, features partial occlusions where the differentiation between a bowl and a vase becomes complex. These examples highlight the need for our model to develop a more nuanced understanding of the scene’s details.
@@ Appendix C Failure Cases, Demo and Future Work
Short Introduction To Demo. We include an introduction video and a demo video in addition to our main paper and supplementary file. The former presents a short introduction to better understand our work, while the latter shows the demo tools of our Open-Vocabulary SAM (please also refer to Fig. 7 for illustration), which can segment and recognize various classes on many scenes. Please refer to our project page: https://www.mmlab-ntu.com/project/ovsam for more details.
@@
Future work. While users can efficiently interact with specific objects using point-and-click or box-dragging techniques, future work will explore using coarse masks or language descriptions as interactive prompts. We aim to continue investigating these promising new directions.

### v2

@@ 4 Experiments
4.1 Main Results
Comparison with Combined Baselines Using Ground Truth. To avoid the influence of other modules, we first demonstrate the recognition ability of our model in Tab. 1. Compared to the simple combined approaches, adding adapters with joint co-training leads to better results. However, the recognition ability is still limited on both COCO and LVIS. Our Open-Vocabulary SAM achieves the best results on both boxes and points as visual prompts. We observe more significant gains on LVIS datasets. We argue that LVIS contains more small objects, which is more challenging than COCO. Our method can solve and lead to over 20% accuracy improvement. Although the segmentation quality is pretty good (about 80 IoU on COCO and LVIS with box prompt), our method still achieves 2% IoU improvements. This indicates the effectiveness of our joint co-training on mask prediction and classification. Compared with boxes as prompts, using points as prompts is more challenging since the location clues of points are much weaker than boxes. However, our approach is still better than combined baselines or them with adapters.
Comparison with Combined Baselines on OV-Detector. In Tab. 2, we adopt a more challenging setting by using the box prediction from the existing open-vocabulary detector to simulate the interactive segmentation process with deviation. We choose the representative Detic [90] as the open-vocabulary detector. Again, our method also achieves the best performance on both COCO and LVIS datasets. In particular, on COCO, compared with previous works [89], our method achieves 3.0 mask mAP improvements with much lower costs.
Comparison with Open-vocabulary Segmentation. In Tab. 3, we compare our method with previous open-vocabulary instance segmentation methods. Open-Vocabulary SAM uses the boxes generated by Mask- RCNN [22] and previous works use the masks of Mask-RCNN. The results show that our Open-Vocabulary SAM performs strongly, especially in the novel classes.
Comparison with SAM on Various Detectors. In Tab. 4, we test the mask prediction quality of our model and original SAM on different detectors. Our method performs better than the original SAM and performs comparably with fine-tuned SAM. It is worth noting that our Open-Vocabulary SAM has much lower computational costs and parameters than SAM.
Comparison on Interactive Segmentation. In Tab. 8 (right), we compare interactive segmentation performance. Notably, our approach excels beyond interactive segmentation and can recognize classes in an open-vocabulary setting.
Visualization Comparison. In Fig. 4, we compare our approach with the feature-crop baseline. Our model shows better performance in classifying small and rare object classification and handling occlusion scenarios.
Model as a Zero Shot Annotation Tool. In addition to COCO and LVIS standard datasets training, following the spirit of SAM, we also scale up our model by training it with more data (Tab. 5). In particular, we adopt more detection data (V3Det [60], Object365 [56]) and classification data (ImageNet22k [11]). Owing to significant costs, we have not conducted comparisons with other baselines for this setting. Rather, we have adapted our method into an interactive annotation tool capable of segmenting and recognizing over 22,000 classes.
4.2 Ablation Studies and Analysis
@@
Setting | IoU | Acc | |||
SAM2CLIP (baseline) | 78.1 | 54.2 | 27.6 | 33.2 | |
+ Cls Token | 81.3 | 79.3 | 29.8 | 34.5 | |
+ Cls Token & CLIP MLP fusion | 80.9 | 78.9 | 29.0 | 33.9 | |
@@
Backbone | IoU | Acc | #FLOPs(G) | #Params (M) |
@@
ViT-L-14 | 38.6 | 14.3 | 2,294 | 441 |
Detailed Design on CLIP2SAM. In Tab. 7, we present extensive design for the CLIP2SAM module. We compare two designs: a simple classification token with cross attention (Cls Token) and a combination of this token with mask pooled CLIP feature (CLS Token & CLIP MLP fusion). These designs work better than the combined baseline shown in the first row. Nonetheless, due to resolution constraints, these variants cannot handle small objects well, as shown in Fig. 4. In contrast, our design improves the performance considerably.
Ablation on Different CLIP Backbones. In Tab. 8, we explore the effect of frozen CLIP visual backbone. We do not add the CLIP2SAM. Motivated by recent works [31, 76, 73, 67], CNN-based CLIPs encapsulate more structural information, which is good for our goal since we have location-sensitive visual prompts. Thus, adopt CNN-based CLIPs, aligned with previous works. As shown in the table, we find ConvNext large achieves the best performance, but we choose RN50x16 since it has comparable performance and better efficiency.
@@ 5 Conclusion
We present Open Vocabulary SAM, a SAM-inspired method for interactive segmentation and recognition. Unlike previous open-vocabulary detection and segmentation methods, our method explores interactive open-vocabulary segmentation for the first time. Given the user’s inputs, such as boxes or points, the proposed approach can interactively segment and label each visual prompt. Compared with the combined baselines and various visual adapters, our proposed CLIP2SAM and SAM2CLIP are both efficient and effective in various settings. Our open vocabulary segmentation is compatible with different detectors, including open-vocabulary detectors and close-set detectors. With more data, our model plays a similar role as SAM, offering an effective annotation tool for both segmentation and instance labeling. In particular, our method can perform large vocabulary segmentation and recognition over 22K classes. We hope our Open-Vocabulary SAM can provide a solid baseline for combining the strengths of different forms of vision foundation models and inspire further research.
@@
Overview. In the supplementary, we provide more details and results to support our main paper. The contents are presented as follows: we first present more details and discussions of our method in Sec. S1. Then, we report more results in Sec. S2. Finally, we present a demo and discuss future work in Sec. S3. Please refer to our GitHub repo at https://github.com/HarborYuan/ovsam for more details about the demo and implementation.
Appendix S1 Method Discussion
S1.1 Comparison with Recent Joint SAM and CLIP Models
Several recent works [6, 83, 59, 20] also explore joint segmentation and recognition as one system. Recognize Anything [83] adopts a paradigm for image tagging. The focus of that work is to build large-scale tagging datasets via automatic text semantic parsing. The SAM model is only an external module. Therefore, it cannot perform well with box or mask-level recognition [83]. SAM-CLIP [59] integrates multi-task learning, continual learning techniques, and teacher-student distillation, where its goal is to build a single model capable of segmentation and classification. However, its classification capability is limited to the image level since it does not provide optimization for small objects and falls short of the original CLIP. Conversely, our Open-Vocabulary SAM fuses knowledge from SAM and CLIP via SAM2CLIP and CLIP2SAM, which keeps the original CLIP and can support interactive segmentation, instance-level classification, and scale-up beyond the frozen CLIP. SSA [6] also builds a system that integrates the knowledge from CLIP and SAM. It involves a semantic voting branch for refining the semantic segmentation results generated by an external semantic segmentation model and refined by SAM. On the contrary, Open-Vocabulary SAM does not require the external segmentor and takes visual prompts rather than coarse semantic segmentation results as input. Moreover, Open-Vocabulary SAM focuses on instance segmentation rather than semantic segmentation. Sambor [20] builds on frozen SAM and CLIP, but adds a SideFormer to fuse the knowledge. The SideFormer serves as an adapter for generating object proposals. Compared to Open-Vocabulary SAM, however, the design keeps two foundation models, which introduces extra computational costs similar to combined baselines. In addition, Sambor does not perform interactive segmentation like Open-Vocabulary SAM.
S1.2 Comparison with Open-Vocabulary Methods
@@
Although our method can recognize objects in an open-vocabulary setting, our work is in a different setting compared to previous open-vocabulary methods. Specifically, we aim to build a SAM-like model with interactive prompts, such as points and boxes. On the contrary, previous open-vocabulary detection methods [90, 18, 78, 79] need to generate region proposals or mask proposals to recall all possible objects in the scene. Thus, most approaches aim to improve the recall of novel objects in region proposal networks or mask proposal networks [79, 9]. Meanwhile, previous open-vocabulary semantic segmentation methods [72, 63, 67, 64] focus on cross-dataset evaluation mode. Our Open-Vocabulary SAM is different from all these works. Our model, following the same spirit of SAM, mainly takes visual proposals for segmentation and recognition. Thus, not only can users deploy our model to take the visual prompts from humans but also deploy our model on various OV-detectors [90, 67] to achieve instance segmentation or class-agnostic detectors to achieve joint segmentation and recognition. Moreover, once scaled up with massive datasets, our method is more practical than previous open-vocabulary works mainly designed for several specific datasets, such as ADE-20k, Pascal-VOC, or LVIS. We provide an interactive demo that can take visual prompts from humans to demonstrate real-world performance.
S1.3 Training and Inference Details
@@
In Fig. S1, we present a more detailed visualization of both training and inference of our Open-Vocabulary SAM. As shown in the figure, only three components are learned during the training, including SAM2CLIP (129M), CLIP2SAM (3.8M), and SAM decoder (4.0M). The remaining parameters are frozen. The total parameters are 304M. After SAM2CLIP distillation, the heavy SAM encoder (637M) is dropped when fine-tuning our model on the detection and classification datasets, which speeds up the CLIP2SAM process. The CLIP2SAM module is co-trained using a diverse mixture of datasets, encompassing detection, segmentation, and classification tasks. Specifically, for segmentation, the training integrates both mask and classification labels, while for detection and classification tasks, the focus is solely on training the classification head. The classification dataset, notable for its extensive range of class categories, coupled with the inclusion of several recent works that feature an extensive vocabulary, significantly enhances the model’s capabilities. After this extensive co-training process, our model exhibits the remarkable ability to recognize and segment over twenty thousand class categories, all within a single, unified framework.
Method | Detectors | mAP | AP50 | AP75 | APS | APM | APL | APr | APn | APf | #Params | FLOPs |
SAM-Huge | Detic (swin-base) | 43.5 | 59.8 | 46.5 | 31.4 | 55.3 | 60.6 | 42.6 | 43.3 | 44.2 | 641M | 3,001G |
Open-Vocabulary SAM | Detic (swin-base) | 42.6 | 59.5 | 45.1 | 29.7 | 53.6 | 61.6 | 42.0 | 42.5 | 42.9 | 304M | 1,180G |
SAM-Huge | ViTDet (Huge) | 44.5 | 61.5 | 47.4 | 34.2 | 56.1 | 60.4 | 34.6 | 45.4 | 47.8 | 641M | 3,001G |
TAP-L [49] | ViTDet (Huge) | 42.6 | – | – | 29.8 | 55.5 | 64.8 | 33.3 | 43.6 | 45.5 | 312M | 1502G |
Open-Vocabulary SAM | ViTDet (Huge) | 44 | 61.7 | 47.1 | 33 | 54.9 | 61.6 | 34.5 | 44.9 | 47.3 | 304M | 1,180G |
Appendix S2 More Experimental Results
S2.1 More Comparison with SAM
In the main paper, we already compare our method with SAM on the COCO dataset (Tab.4), which shows that Open-Vocabulary SAM has good performance compared to the original SAM. In Tab. S2, we further compare our method with SAM and a recent SAM-like method TAP [49]. The results of TAP-L [49] are from the original paper. Regarding the mask quality, we notice that Open-Vocabulary SAM has comparable or slightly worse results compared to the original SAM. However, we want to note that Open-Vocabulary SAM has a much lower computational cost. When comparing with recently proposed TAP [49], Open-Vocabulary SAM has better performance and lower computational cost.
S2.2 More Visualization Comparison
In Fig. S2, we compare our method with the feature-crop baseline with visualization results. We noticed that although the recognition capability of the feature-crop baseline is significantly poor compared to Open-Vocabulary SAM, especially on the small objects, the errors are relatively traceable. We noticed that the feature-crop baseline can extract useful information from features, but it may be confused with adjacent objects. We speculate that this may be because it only performs image-level training and lacks object-level recognition capabilities. In contrast, Open-Vocabulary SAM has great advantages at object-level recognition, especially in the recognition of small objects.
S2.3 Segment and Recognize Everything
In Fig. S3, we demonstrate demonstrate the capabilities of Open-Vocabulary SAM on segment and recognize everything. With the box prompts as input, our method can segment every identifiable object within a given image, showcasing its versatility and effectiveness in handling diverse objects. Based on the results presented in Fig. S3, Open-Vocabulary SAM not only excels in segmenting a wide range of objects but also in recognizing them, which shows its potential as a powerful tool for comprehensive image analysis.
Appendix S3 Failure Cases, Demo and Future Work
S3.1 Failure Cases
In Figure S4, we present a series of instances where our Open-Vocabulary SAM does not perform well. First, Open-Vocabulary SAM struggles to differentiate between classes that are subtly distinct from one another, such as mistaking "Truck" for "Minivan". Moreover, some challenging cases, such as extreme occlusion, also confuse Open-Vocabulary SAM. For example, Open-Vocabulary SAM incorrectly classified “Vase” as “Bowl” in the bottom right example. Besides these challenging scenarios, another notable and interesting issue is the model’s confusion with overlapping objects. It misclassified the “Toy” to the “Toothbrush” (the third example in the figure), where the two objects have significant overlapping. These examples highlight areas where further research could be beneficial in enhancing the performance of Open-Vocabulary SAM.
S3.2 Short Introduction To Demo
@@
We present an online demo in addition to our paper. It shows the functionality of our Open-Vocabulary SAM (please also refer to Fig. S5 for illustration), which can segment and recognize various classes on many scenes. Please refer to our GitHub repo at https://github.com/HarborYuan/ovsam for more details.
S3.3 Future Work
@@
While users can efficiently interact with specific objects using point-and-click or box-dragging techniques, future work will explore using coarse masks or language descriptions as interactive prompts. We aim to continue investigating these promising new directions. Also, our work mainly focuses on foreground categories (e.g., person), and future work may consider introducing background categories (e.g., sky) for comprehensive segmentation and recognition.
Acknowledgements. This study is supported under the RIE2020 Industry Alignment Fund-Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contributions from the industry partner(s). The project is also supported by Singapore MOE AcRF Tier 1 (RG16/21) and the National Key R&D Program of China (No. 2022ZD0161600).